OpenStudy (anonymous):

What is the y-intercept of y = -8(5)x? -8 5 1 8

OpenStudy (anonymous):

100000

OpenStudy (cwrw238):

y intercept is the value of y when x = 0

mathslover (mathslover):

r u sure that the question is y = -8(5)x ? since to find y intercept we have to put x = 0 , y = 0 that is not in the options

OpenStudy (anonymous):

I think 0= x??

mathslover (mathslover):

yes @rachelbreezy to find y intercept we put x = 0 and vice versa .. right

OpenStudy (anonymous):

Im confused

Parth (parthkohli):

I think he meant \(y = -8(5^x)\) Then set x = 0 \( \color{Black}{\Rightarrow y = -8(5^0) }\) \( \color{Black}{\Rightarrow y = -8 \cdot 1}\) \( \color{Black}{\Rightarrow y = -8}\)

Parth (parthkohli):

she*

mathslover (mathslover):

hmn then it is ok !

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

no